# Simon Briend

### v2

@@ A.5 Proof of the minimax lower bound in the PA model
$\geq$ $\frac{1}{4}\left(1-\frac{5}{n}\right)^{5}~{}.$
@@ A.6 Descendant ordering in the PA model
  1. 1.
    $\forall v\in[n]\backslash\{1\to c\},\ \psi_{T}(v)=\psi^{\prime}_{T}(v)~{};$
  2. 2.
@@
    $\widehat{\sigma}_{J}(1)\leq\frac{C}{\epsilon^{2}}\exp\left(\sqrt{C\log\left(\frac{1}{\epsilon}\right)}\right)~{}.$
@@
The first part of the proof is identical to the proof of Lemma 2. First, we use Theorem 6 of Wagner and Durant [31], which extend the result of Moon [23] from uniform random recursive trees to preferential attachment trees. Using their result, we obtain that
$\mathbb{P}\left\{c\geq k\right\}\leq\sum_{j=k}^{\infty}\frac{(-\log(2)/2)^{j}}{j!}~{},$
@@
$\widehat{\sigma}_{J}(1)\leq\frac{C}{\epsilon^{2}}\exp\left(\sqrt{C\log\left(\frac{1}{\epsilon}\right)}\right)~{},$
@@
$R_{\alpha}(\widehat{\sigma}_{J})\leq R_{\alpha}(\widehat{\sigma}^{\prime})+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{2}(n)\sqrt{n}~{}.$
@@
$\leq\frac{1}{2}\mathbb{E}\left[D^{2}\right]+\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]~{}.$
@@
$\frac{1}{2}\mathbb{E}\left[D^{2}\right]\leq K^{2}~{}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\left(\mathbb{E}\left[\widehat{\sigma}_{J}(1)^{1+\gamma}\right]\right)^{\frac{1}{1+\gamma}}~{}.$ (16)
@@
$\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\leq C\frac{1}{1+\gamma}{\gamma}~{},$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq f(\epsilon)\right\}\leq\epsilon~{},$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq k\right\}\leq\frac{C}{\sqrt{k}}\exp\left(\sqrt{C\log(k)}\right)~{}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq C\log^{2}(n)\sqrt{n}~{},$
@@ A.7 Performance of Jordan ordering in the PA model
Similarly to the urrt case, in the pa model, the number of descendants of a vertex is distributed as a Pólya urn. This well-know fact is easily seen since in the pa model, sampling a vertex with a probability proportional to its degree is the same as sampling an edge uniformly at random and picking one of its endpoints at random. In turn, it is the same as picking a half edge uniformly at random. Therefore, the resulting Pólya urn has slightly different initial conditions than in the urrt. Such Pólya urns are well understood. In particular, by Mahmoud [22, Section 3.2], for a vertex $i\in[n]$, the distribution of ${\rm de}(i)$ is given by
$\mathbb{P}\left\{{\rm de}(i)=k\right\}=\frac{\left(13\cdots(2k-1)\right)\left((2i-3)(2i-1)\cdots(2n-2k-5)\right)}{(2i-2)2i\cdots(2n-4)}\binom{n-i}{k}~{}.$
@@
$\cdot\underbrace{\frac{(n-i)!/(n-i-k)!}{(2n-2k-3)(2n-2k-1)\cdots(2n-5)}}_{\stackrel{{\scriptstyle\rm def}}{{=}}C}~{}.$
@@
$A=\frac{1}{k}\prod_{j=1}^{k-1}\frac{2j+1}{j}=\frac{2^{k-1}}{k}\prod_{j=1}^{k-1}\left(1+\frac{1}{2j}\right)~{}.$
@@
$\prod_{j=1}^{k-1}\left(1+\frac{1}{2j}\right)=\exp\left(\sum_{j=1}^{k-1}\log\left(1+\frac{1}{2j}\right)\right)\leq\exp\left(\sum_{j=1}^{k-1}\frac{1}{2j}\right)\leq\sqrt{k}~{},$
@@
$=\exp\left(\sum_{j=i}^{n-1}\log\left(1-\frac{1}{2j-2}\right)\right)\leq\exp\left(-\sum_{j=i}^{n-1}\frac{1}{2j-2}\right)\leq 2\sqrt{\frac{i}{n}}~{}$
@@
$C=\prod_{j=n-k-2}^{n-3}\frac{j-i-3}{2j+1}=\frac{1}{2^{k-1}}\prod_{j=n-k-2}^{n-3}\left(1-\frac{i+2.5}{j+0.5}\right)\leq\frac{1}{2^{k-2}}\left(1-\frac{k}{n}\right)^{i}~{}.$
@@
$\mathbb{P}\left\{{\rm de}(i)=k\right\}\leq 4\sqrt{\frac{i}{kn}}\left(1-\frac{k}{n}\right)^{i}~{}.$
@@
$\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau\right\}\leq 10\sqrt{i\tau}~{}.$ (18)
@@
$\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau\right\}\leq 10\sqrt{\frac{i}{2j}}\sqrt{1+\log(j/i)}~{},$
@@
$\leq 8\sqrt{2}\sqrt{\frac{i}{j}}~{}.$
@@
$\mathbb{P}\left\{{\rm de}(i)\leq{\rm de}(j)\right\}\leq\left(8\sqrt{2}+\frac{10}{\sqrt{2}}\sqrt{1+\log(j/i)}\right)\sqrt{\frac{i}{j}}~{}.$
@@
$\leq\frac{2}{2-\alpha}n^{2-\alpha}+\sum_{j=1}^{n}\left(\frac{1}{\sqrt{j}}\sum_{i=1}^{j-1}\left(8\sqrt{2}+\frac{10}{\sqrt{2}}+\frac{10}{\sqrt{2}}\log(j/i)\right)i^{1/2-\alpha}\right)~{}.$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq\left(\frac{2}{2-\alpha}+\frac{8\sqrt{2}+\frac{10}{\sqrt{2}}}{(3/2-\alpha)(2-\alpha)}+\frac{20}{\sqrt{2}(2-\alpha)(3/2-\alpha)^{2}}\right)n^{2-\alpha}~{},$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq Cn^{3/2}~{},$

### v3

@@ A.5 Proof of the minimax lower bound in the PA model
$\geq$ $\frac{1}{4}\left(1-\frac{5}{n}\right)^{5}.$
@@ A.6 Descendant ordering in the PA model
  • $\forall v\in[n]\backslash\{1\to c\},\ \psi_{T}(v)=\psi^{\prime}_{T}(v);$
@@
    $\widehat{\sigma}_{J}(1)\leq\frac{C}{\epsilon^{2}}\exp\left(\sqrt{C\log\left(\frac{1}{\epsilon}\right)}\right).$
@@
The first part of the proof is identical to the proof of Lemma 2. First, we use Theorem 6 of [13], which extend the result of Moon [24] from uniform random recursive trees to preferential attachment trees. Using their result, we obtain that
$\mathbb{P}\left\{c\geq k\right\}\leq\sum_{j=k}^{\infty}\frac{(-\log(2)/2)^{j}}{j!},$
@@
$\widehat{\sigma}_{J}(1)\leq\frac{C}{\epsilon^{2}}\exp\left(\sqrt{C\log\left(\frac{1}{\epsilon}\right)}\right),$
@@
$R_{\alpha}(\widehat{\sigma}_{J})\leq R_{\alpha}(\widehat{\sigma}^{\prime})+K\sum_{i=1}^{n}\frac{1}{i^{\alpha}}+C\log^{2}(n)\sqrt{n}.$
@@
$\leq\frac{1}{2}\mathbb{E}\left[D^{2}\right]+\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right].$
@@
$\frac{1}{2}\mathbb{E}\left[D^{2}\right]\leq K^{2}.$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\left(\mathbb{E}\left[\widehat{\sigma}_{J}(1)^{1+\gamma}\right]\right)^{\frac{1}{1+\gamma}}.$ (16)
@@
$\left(\mathbb{E}\left[D^{\frac{1+\gamma}{\gamma}}\right]\right)^{\frac{\gamma}{1+\gamma}}\leq C\frac{1}{1+\gamma}{\gamma},$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq f(\epsilon)\right\}\leq\epsilon,$
@@
$\mathbb{P}\left\{\widehat{\sigma}_{J}(1)\geq k\right\}\leq\frac{C}{\sqrt{k}}\exp\left(\sqrt{C\log(k)}\right).$
@@
$\mathbb{E}\left[D\widehat{\sigma}_{J}(1)\right]\leq C\log^{2}(n)\sqrt{n},$
@@ A.7 Performance of Jordan ordering in the PA model
Similarly to the urrt case, in the pa model, the number of descendants of a vertex is distributed as a Pólya urn. This well-know fact is easily seen since in the pa model, sampling a vertex with a probability proportional to its degree is the same as sampling an edge uniformly at random and picking one of its endpoints at random. In turn, it is the same as picking a half edge uniformly at random. Therefore, the resulting Pólya urn has slightly different initial conditions than in the urrt. Such Pólya urns are well understood. In particular, by Mahmoud [23, Section 3.2], for a vertex $i\in[n]$, the distribution of ${\rm de}(i)$ is given by
$\mathbb{P}\left\{{\rm de}(i)=k\right\}=\frac{\left(13\cdots(2k-1)\right)\left((2i-3)(2i-1)\cdots(2n-2k-5)\right)}{(2i-2)2i\cdots(2n-4)}\binom{n-i}{k}.$
@@
$\cdot\underbrace{\frac{(n-i)!/(n-i-k)!}{(2n-2k-3)(2n-2k-1)\cdots(2n-5)}}_{\stackrel{{\scriptstyle\rm def}}{{=}}C}.$
@@
$A=\frac{1}{k}\prod_{j=1}^{k-1}\frac{2j+1}{j}=\frac{2^{k-1}}{k}\prod_{j=1}^{k-1}\left(1+\frac{1}{2j}\right).$
@@
$\prod_{j=1}^{k-1}\left(1+\frac{1}{2j}\right)=\exp\left(\sum_{j=1}^{k-1}\log\left(1+\frac{1}{2j}\right)\right)\leq\exp\left(\sum_{j=1}^{k-1}\frac{1}{2j}\right)\leq\sqrt{k},$
@@
$=\exp\left(\sum_{j=i}^{n-1}\log\left(1-\frac{1}{2j-2}\right)\right)\leq\exp\left(-\sum_{j=i}^{n-1}\frac{1}{2j-2}\right)\leq 2\sqrt{\frac{i}{n}}.$
@@
$C=\prod_{j=n-k-2}^{n-3}\frac{j-i-3}{2j+1}=\frac{1}{2^{k-1}}\prod_{j=n-k-2}^{n-3}\left(1-\frac{i+2.5}{j+0.5}\right)\leq\frac{1}{2^{k-2}}\left(1-\frac{k}{n}\right)^{i}.$
@@
$\mathbb{P}\left\{{\rm de}(i)=k\right\}\leq 4\sqrt{\frac{i}{kn}}\left(1-\frac{k}{n}\right)^{i}.$
@@
$\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau\right\}\leq 10\sqrt{i\tau}.$ (18)
@@
$\mathbb{P}\left\{\frac{{\rm de}(i)}{n}\leq\tau\right\}\leq 10\sqrt{\frac{i}{2j}}\sqrt{1+\log(j/i)},$
@@
$\leq 8\sqrt{2}\sqrt{\frac{i}{j}}.$
@@
$\mathbb{P}\left\{{\rm de}(i)\leq{\rm de}(j)\right\}\leq\left(8\sqrt{2}+\frac{10}{\sqrt{2}}\sqrt{1+\log(j/i)}\right)\sqrt{\frac{i}{j}}.$
@@
$\leq\frac{2}{2-\alpha}n^{2-\alpha}+\sum_{j=1}^{n}\left(\frac{1}{\sqrt{j}}\sum_{i=1}^{j-1}\left(8\sqrt{2}+\frac{10}{\sqrt{2}}+\frac{10}{\sqrt{2}}\log(j/i)\right)i^{1/2-\alpha}\right).$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq\left(\frac{2}{2-\alpha}+\frac{8\sqrt{2}+\frac{10}{\sqrt{2}}}{(3/2-\alpha)(2-\alpha)}+\frac{20}{\sqrt{2}(2-\alpha)(3/2-\alpha)^{2}}\right)n^{2-\alpha},$
@@
$\sum_{i=1}^{n}\mathbb{E}\left[\frac{|\widehat{\sigma}^{\prime}(i)-i|}{i^{\alpha}}\right]\leq Cn^{3/2},$